Please complete the problems
This quiz is due on Monday, March 12, at the beginning of class. . You must use pencil on this quiz. No exceptions! . You must explain your reasoning or show a calculation on all problems for full credit. Identify the given random variable as being discrete or continuous. 1) The number of text messages between Ken and his girlfriend on St. Patricks Day Determine whether the following is a probability distribution. If not, identify the requirement that is not satisfied 2) It is reported that in Warren, MI 60% of adults have a college degree. The accompanying table describes the probability distribution for the number of adults (among 4 randomly selected adults) who have a college degree. P(x) 0 0.0256 1 0.1536 2 0.3456 3 0.3456 40.1296 Find the mean of the given probability distribution. 3) The Detroit police department reports that the probabilities that 0, 1, 2, and 3 burglaries will be reported in given day are 0.45,0.45, 0.09, and 0.01, respectively. John randomly selects 14 newborm babies and counts the number of boys selected, x. The probabiliti corresponding to the 14 possible values of x are summarized in the given table. Answer the question using the tab Probabilities of Boys 0 0.0005 0.122 10 0.061 | |0001 | 6 |0.183 | 11 |0.022 2 |0006| 7 |0.209| 12 |0.006 3 0.022 0.183 13 0.001 4 0.0619 0.122 14 0.000 4) Find the probability of selecting 12 or more boys B-1
Answer #1

1) Discrete Variable as

A discrete variable is a variable whose value is obtained by counting.

2)

yes it is a probability distribution

as the sum of all the probabilities is 1

3)

Mean= x*P(x)= 0*0.45 +1*0.45 +2*0.09 +3*0.01= 0+0.45+0.18+0.03= 0.66

4)

P(x>=12)=0.006+0.001+0.000= 0.007